Sol Summary Report – February 21th

Crew 275 Sol Summary Report 21Feb2023

Sol: 9

Summary Title: Aluminum foil and zip ties are an astronaut’s best friends

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Even if the food stocks are quickly decreasing, which makes the crew quite anxious about the fact that they will have to limit their food intake more than they already do, several successes have managed to keep our spirits high.

Another EVA dedicated to the atmospheric instruments was successfully conducted today, with the installation of the LOAC, measuring particles in the air. Following the failure of some tests during the previous EVAs, the antenna from Mega-Ares was also retrieved so that we can fix it from within the Hab. A smart fix based on aluminum foil and aluminum tape has indeed been found and all the atmospheric instruments should be fully functional very soon.

In the meantime, the photogrammetry team has finished generating 3D and 2D maps of the terrain explored yesterday at North Ridge. They started training the crewmembers taking part in the next EVAs to explain the tasks they will have to perform onsite, so that it is possible to evaluate the added value of a 3D map over a 2D map.

We started a human factor experiment whose purpose is to evaluate the impact of the architecture of an extraterrestrial station on the overall performance of the crew. To do so, we conducted very demanding cognitive tests at different locations through the campus: the Upper Deck of the Hab, the Science Dome, and the Green Hab. This experiment includes also other protocols that will be implemented progressively.

In the evening, because of the wind, the external door of the main airlock was blown open and we had to secure it with a rope. For safety’s sake, the door of the engineering air lock was also secured with a screwdriver.

Look Ahead Plan:

With the 2D and 3D maps of the terrain explored yesterday fully rendered, we are planning to send two teams to the exact same location to check the added value of the 3D map to ease the ability of astronauts to identify some specific features onsite. We thus plan to send an EVA crew to North Ridge with a 2D map tomorrow morning (Sol 10) and another one with a 3D map on Thursday morning (Sol 11).

As explained yesterday, we would like to perform another EVA tomorrow (Sol 10) in the afternoon to put back the antenna of Mega-Ares back at its initial place. The antenna has indeed been fixed today.

With the LOAC instrument now deployed, it should also be noted that we would like to have the possibility to perform an “emergency EVA” in case too strong precipitations are announced, or at least to take the opportunity of another EVA to retrieve it. Indeed, this instrument is not only very expensive, but also very sensitive to precipitations and could be damaged if it rains or snows too much.

Anomalies in work: None.

Weather: Cloudy and windy, low -4°C/24°F, high 8°C/47°F.

Crew Physical Status: Optimal.

EVA: The team of EVA#7 successfully deployed the LOAC instrument and retrieved the Mega-Ares antenna to repair it inside the hab.

Reports to be filed: Sol Summary, EVA Request, EVA Report, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site